Required information The following information applies to the questions displayed below] Tyrell Co. entered into the followinRequired information [The following information applies to the questions displayed below.] Tyrell Co. entered into the followRequired information [The following information applies to the questions displayed below.] Tyrell Co. entered into the follow! Required information [The following information applies to the questions displayed below. Tyrell Co. entered into the folloо Required information [The following information applies to the questions displayed below.] Tyrell Co. entered into the foll
Answer #1
1) Maturity date
locust NBR fargo
date of the note 19-May 8-Jul 28-Nov
term of note 90 120 60
maturity date 17-Aug 5-Nov 27-Jan
2) interest due at maturity
principal * Rate * time = interest
locust 35,000 * 8% * 90/360 = 700
NBR 60,000 * 11% * 120/360 = 2200
Fargo 27,000 * 7% * 60/360 = 315
3) Amount in adjusting entry
Fargo Bank
principal * Rate * time = interest
interest to be acccrued in 2016 27,000 * 7% * 33/360 = 173
4) interest expense to be recorded in 2017
principal * Rate * time = interest
interest to recorded in 2018 27,000 * 7% * 27/360 = 142
Journal entries
Date Accounting titles & Explanations Debit Credit
2016
20-Apr inventory 37,500
Accounts payable 37,500
19-May Accounts payable 37,500
cash 2,500
notes payable 35,000
8-Jul Cash 60,000
notes payable 60,000
17-Aug notes payable 35,000
interest expense 700
cash 35,700
5-Nov notes payable 60,000
interest expense 2,200
cash 62,200
28-Nov Cash 27,000
notes payable 27,000
31-Dec interest expense 173
interest payable 173
2017
27-Jan notes payable 27,000
interest payable 173
interest expense 142
cash 27,315